# Counting Salem numbers arising from arithmetic hyperbolic orbifolds

### v1

@@ 1. Introduction
The condition $f(1)f(-1)\equiv-D$ has applications to the study of arithmetic hyperbolic orbifolds. Indeed, Bayer-Fluckiger [BF15] stated this condition as a necessary condition for the existence of an isometry in a quadratic space with characteristic polynomial $f$. It is known that Salem numbers are closely related to the lengths of geodesics in arithmetic hyperbolic orbifolds. Recent work of Emery, Ratcliffe, and Tschantz made this relationship more explicit.
@@
This question was addressed in [BLMT22] in the case where $m=2$, $n=3$, and $\Gamma$ is a Bianchi group. We obtain the following theorem, which bounds the proportion of Salem numbers of degree up to $n+1$ in the commensurability class of classical arithmetic hyperbolic lattices in odd dimensions $n$. Note that this theorem applies to all non-cocompact arithmetic lattices in odd dimensions but also to some cocompact lattices in dimension 3.
@@
Let $q$ be a rational quadratic form with signature $(n,1)$ with $n$ odd. The number of Salem numbers $\lambda\leq Q$ realized in the commensurability class of $\mathrm{SO}(q,\mathbb{Z})$ is bounded above by
@@
Let $n$ be odd and let $\Gamma$ be a classical arithmetic lattice in $\mathrm{Isom}(\mathbb{H}^{n})$ commensurable with some $\mathrm{O}(q,\mathbb{Z})$ for an admissible quadratic form $q$ over $\mathbb{Q}$. Suppose $q$ has reduced determinant D-D. Let $m=\frac{n+1}{2}$. The number of distinct lengths $\leq L$ of hyperbolic isometries which are elements of classical arithmetic lattices in the commensurability of $\Gamma$ is bounded above by
@@
Let $n$ be odd and let $\Gamma$ be an arithmetic lattice in $\mathrm{Isom}(\mathbb{H}^{n})$ commensurable with some $\mathrm{O}(q,\mathbb{Z})$ for an admissible quadratic form $q$ over $\mathbb{Q}$. Suppose $q$ has reduced determinant D-D. Let $m=\frac{n+1}{2}$. The number of distinct lengths $\leq L$ of hyperbolic isometries which are elements of classical arithmetic lattices in the commensurability of $\Gamma$ is bounded above by
$\frac{4\kappa(m,D)}{\pi\sqrt{D}}e^{2(m-1)L}L+O(e^{2(m-1)L})$
@@
We apply our methods to improve lower bounds for the strong exponential growth of average of multiplicities in the geodesic length spectrum of non-compact arithmetic orbifolds. This was initiated in [BLMT22] in the even case and in [Gre24] in the odd case.
@@ Acknowledgements
The first author was partially supported by NSF grant DMS-2243188. The second author was partially supported by Universal CNPq Grant 408834/2023-4. The third author was supported by the SECIHTI Grant 169113, Estancias Posdoctorales por México 2023-Modalidad Académica, at CIMAT.
@@ 2.4. Salem numbers and lengths of geodesics
which can be viewed as the geometric mean of $|p(z)|$ over all values of $z$ lying on the unit circle. One can immediately observe that $M(p)=1$ precisely when $p(x)=x^{k}\displaystyle\prod_{i}\Phi_{n_{i}}(x)$. There is a lower bound on the Mahler measure in terms of the coefficients (see [Mah62, §2])
@@ 3. Counting Diophantine solutions
Let $\mathcal{C}\in\mathbb{R}^{2}$ be a bounded, convex, centrally-symmetric region. As a bounded set, there exists an $R\in\mathbb{R}_{>0}$ such that any point $(x,y)\in\mathcal{C}$ satisfies $R\leq x\leq R$ and $R\leq y\leq R$.
@@
$\alpha\log(\alpha R)+\alpha\gamma+O\left(\frac{1}{R}\right)-\alpha^{2}\mathrm{vol}(\mathcal{C})\sum_{m=\lceil\alpha R\rceil}^{\infty}\frac{\mu(m)}{m^{2}}$
@@
Collecting the information from all four cases, we see that the contribution from even values of $D$ yields $2^{t}\frac{3X}{4\pi\sqrt{D}}+O(\sqrt{X}\log X)$ solutions to the Diophantine equation $A^{2}+DB^{2}=C^{2}$. On the other hand, the contribution from odd values of $D$ yields $2^{t}\frac{X}{2\pi\sqrt{D}}+O(\sqrt{X}\log X)$ solutions. Observe that the above arguments only count positive solutions. Allowing for all $2^{3}$ combinations of $\pm$ solutions yields the result in Theorem 3.1. ∎
@@ 4. Counting Salem numbers
With $A,C,Q$ defined as above, we have $C\leq 2^{2m-1}Q$ and
$A\leq(2^{2m-1}-2)Q+2$.
@@
$C=\sum_{k\text{ odd}}c_{k}\leq\sum_{k\text{ odd}}|c_{k}|\leq\sum_{k\text{ odd}}\binom{2m}{k}Q=2^{2m-1}Q.$
@@
Theorem 3.1 allows us to count the number of integer solutions for the triple $(A,B,C)$ satisfying $A^{2}+DB^{2}=C^{2}$ with $|C|\leq 2^{2m-1}Q$. However, we are interested in counting the integer solutions only for the tuple $(A,C)$, so we ignore the sign of $B$ by dividing by $2$. Moreover, we only want to count the subset of solutions where $C$ is even whenever $m$ is even and $A$ is even whenever $m$ is odd (see the proof of Lemma 4.4), so we again divide by $2$. This results in the following count for admissible integral solutions $(A,C)$:
@@
$\mathcal{F}^{red}_{m,D}(Q)\leq\sum_{i=1}^{m-1}\iota_{0}(i)\cdot\#\mathcal{F}^{\prime}_{m-i}(Q)\leq\sum_{i=1}^{m-1,D}\iota_{0}(i)\cdot\frac{\kappa(i,D)}{\pi\sqrt{D}}Q^{m-i-1}\log Q\leq\iota(m,D)Q^{m-2}\log Q$
@@ 5. Counting lengths of closed geodesics in commensurability classes
Let $r$ denote the rank of $q$, which is also the degree of $p_{T}\in\mathbb{Z}[x]$. The matrix powers $T,T^{2},\dots,T^{n-1}$ all have entries in $\mathbb{Q}$. Let $\ell$ denote the least common multiple of the denominators appearing in the entries of these n1n-1 rational matrices. By the Cayley-Hamilton Theorem, $T$ satisfies its characteristic polynomial, so for any $k\geq 0$, we have
$T^{k+r}=-\left(T^{k}+c_{1}T^{k+1}+\cdots+c_{r-1}T^{k+r-1}\right).$
The subgroup of $\mathbb{Q}^{n}$ generated by the $T^{k}\mathbb{Z}^{r}$ for all $k\geq 0$ lies between the subgroups $\mathbb{Z}^{r}$ and $(\frac{1}{\ell}\mathbb{Z})^{r}$ and is thus itself a free abelian group of rank $r$. We can therefore choose a change of basis matrix $g\in\mathrm{GL}(n+1,\mathbb{Q})$ such that $g^{-1}Tg$ has entries in $\mathbb{Z}$.
@@
The Salem numbers which are realizable in the commensurability class of $\mathrm{SO}(q,\mathbb{Z})$ have even degree $2m\leq n+1$. Let $\mathcal{R}_{q}(Q)$ denote the set of realizable Salem numbers with $\lambda<Q$. By [MA13], the reduced determinant of a rational quadratic form of even rank is a commensurability invariant. By [BF15, Corollary 9.3], a necessary condition for the existence of an isometry in $\mathrm{O}(q,\mathbb{Q})$ with minimal polynomial $f$ is that $\det(q)\equiv f(1)f(-1)$ in $\mathbb{Q}^{\times}/\mathbb{Q}^{\times 2}$. Since $f(0)=1$, such an isometry will be in $\mathrm{SO}(q,\mathbb{Q})$, and hence by Lemma 5.1, it will be in the commensurability class of $\mathrm{SO}(q,\mathbb{Z})$. Therefore we have
@@
We can now deduce C.
Proof of C.
Recall that the length of a hyperbolic element $\gamma\in\Gamma$ satisfies $\ell(\gamma)=\log\lambda$ where $\lambda$ is the unique real eigenvalue of $\gamma$ such that $\lambda>1$. Then $\ell(\gamma)\leq L$ if and only $\lambda\leq e^{L}$. The result follows from B by replacing $Q$ with $e^{L}$. ∎
@@ 6. Bounds on multiplicity of the length spectrum
For arithmetic hyperbolic orbifolds, a lower bound for the average of the multiplicities can be obtained by counting Salem numbers. This study was initiated in [BLMT22] for even $n$, and in [Gre24] for odd $n$. Both of these papers used results about the so-called square-rootable Salem numbers. Let

### v2

@@ 1. Introduction
The condition $f(1)f(-1)\equiv-D$ has applications to the study of arithmetic hyperbolic orbifolds. Indeed, Bayer-Fluckiger [BAY15] stated this condition as a necessary condition for the existence of an isometry in a quadratic space with characteristic polynomial $f$. It is known that Salem numbers are closely related to the lengths of geodesics in arithmetic hyperbolic orbifolds. Recent work of Emery, Ratcliffe, and Tschantz made this relationship more explicit.
@@
This question was addressed in [BLM+22] in the case where $m=2$, $n=3$, and $\Gamma$ is a Bianchi group. We obtain the following theorem, which bounds the proportion of Salem numbers of degree up to $n+1$ in the commensurability class of classical arithmetic hyperbolic lattices in odd dimensions $n$. Note that this theorem applies to all non-cocompact arithmetic lattices in odd dimensions but also to some cocompact lattices in dimension 3.
@@
Let $q$ be a rational quadratic form with signature $(n,1)$ with $n$ odd and reduced determinant D-D. The number of Salem numbers $\lambda\leq Q$ realized in the commensurability class of $\mathrm{SO}(q,\mathbb{Z})$ is bounded above by
@@
Let $n$ be odd and let $\Gamma$ be a classical arithmetic lattice in $\mathrm{Isom}(\mathbb{H}^{n})$ commensurable with some $\mathrm{O}(q,\mathbb{Z})$ for an admissible quadratic form $q$ over $\mathbb{Q}$. Suppose $q$ has reduced determinant D-D. Let $m=\frac{n+1}{2}$. The number of distinct lengths $\leq L$ of hyperbolic isometries which are elements of classical arithmetic lattices in the commensurability class of $\Gamma$ is bounded above by
@@
Let $n$ be odd and let $\Gamma$ be an arithmetic lattice in $\mathrm{Isom}(\mathbb{H}^{n})$ commensurable with some $\mathrm{O}(q,\mathbb{Z})$ for an admissible quadratic form $q$ over $\mathbb{Q}$. Suppose $q$ has reduced determinant D-D. Let $m=\frac{n+1}{2}$. The number of distinct lengths $\leq L$ of hyperbolic isometries which are elements of classical arithmetic lattices in the commensurability class of $\Gamma$ is bounded above by
$\frac{\kappa(m,D)}{\pi\sqrt{D}}e^{2(m-1)L}L+O(e^{2(m-1)L})$
@@
The study of multiplicities in the length spectrum of arithmetic hyperbolic manifolds is an active area of research. The exponential growth of multiplicity averages for surfaces follows from work by Luo and Sarnak [LS94]. A similar result has recently been obtained for semi-arithmetic hyperbolic surfaces [BCD+26]. For $n\geq 3$, this was initiated in [BLM+22] for the even case and in [GRE24] for the odd case. We apply our methods to improve lower bounds for the strong exponential growth of average of multiplicities in the geodesic length spectrum of non-compact arithmetic orbifolds.
@@ Acknowledgements
The first author was partially supported by NSF grants DMS-2243188 and DMS-2441034. The second author was partially supported by Universal CNPq Grant 408834/2023-4. The third author was supported by the SECIHTI Grant 169113, Estancias Posdoctorales por México 2023-Modalidad Académica, at CIMAT. The fourth author was partially supported by the Max Planck Institute for Mathematics and the Centre de Recherches Mathématiques during her sabbatical.
@@ 2.4. Salem numbers and lengths of geodesics
which can be viewed as the geometric mean of $|p(z)|$ over all values of $z$ lying on the unit circle. One can immediately observe that $M(p)=1$ precisely when $p(x)=x^{k}\displaystyle\prod_{i}\Phi_{n_{i}}(x)$, where $\Phi_{n}$ denotes the $n$-th cyclotomic polynomial. There is a lower bound on the Mahler measure in terms of the coefficients (see [MAH62, §2])
@@ 3. Counting Diophantine solutions
Let $D$ be a square-free positive integer and $t$ be the number of distinct prime factors of $D$. Denote the number of integer solutions to $A^{2}+DB^{2}=C^{2}$ with $\gcd(A,C)=1$ and $|C|\leq X$ by
$P_{D}(X)=\#\{(A,B,C)\in\mathbb{Z}^{3}\mid A^{2}+DB^{2}=C^{2},\,\gcd(A,C)=1,\text{ and }|C|\leq X\}.$
@@
Let $\mathcal{C}\subset\mathbb{R}^{2}$ be a bounded, convex, centrally-symmetric region. As a bounded set, there exists an $R\in\mathbb{R}_{>0}$ such that any point $(x,y)\in\mathcal{C}$ satisfies $R\leq x\leq R$ and $R\leq y\leq R$.
@@
$O\left(\alpha\log(\alpha R)+\alpha\gamma+\frac{1}{R}\right)-\alpha^{2}\mathrm{vol}(\mathcal{C})\sum_{m=\lceil\alpha R\rceil}^{\infty}\frac{\mu(m)}{m^{2}}$
@@
Collecting the information from all four cases, we see that the contribution from even values of $D$ yields $2^{t}\frac{3X}{4\pi\sqrt{D}}+O(\sqrt{X}\log X)$ solutions to the Diophantine equation $A^{2}+DB^{2}=C^{2}$. On the other hand, the contribution from odd values of $D$ yields $2^{t}\frac{X}{\pi\sqrt{D}}+O(\sqrt{X}\log X)$ solutions. Observe that the above arguments only count positive solutions. Allowing for all $2^{3}$ combinations of $\pm$ solutions yields the result in Theorem 3.1. ∎
@@ 4. Counting Salem numbers
With $A,C,Q$ defined as above, we have $|C|\leq 2^{2m-1}Q$ and
$|A|\leq(2^{2m-1}-2)Q+2$.
@@
$|C|=|\sum_{k\text{ odd}}c_{k}|\leq\sum_{k\text{ odd}}|c_{k}|\leq\sum_{k\text{ odd}}\binom{2m}{k}Q=2^{2m-1}Q.$
@@
Theorem 3.1 allows us to count the number of integer solutions for the triple $(A,B,C)$ satisfying $A^{2}+DB^{2}=C^{2}$ with $|C|\leq 2^{2m-1}Q$. However, we are interested in counting the integer solutions only for the tuple $(A,C)$, so we ignore the sign of $B$ by dividing by $2$. Moreover, we only want to count the subset of solutions where $C$ is even whenever $m$ is even and $A$ is even whenever $m$ is odd (see the proof of Lemma 4.4), so we again divide by $2$. This results in the following count for admissible integral solutions $(A,C)$ with $\gcd(A,C)=1$:
@@
However, we are interested in all integral solutions and not only these primitive solutions. To compute this quantity, we observe that:
$\#\{(A,C)\in\mathbb{Z}^{2}:A^{2}+DB^{2}=C^{2}\ \mathrm{with}\ |C|\leq Q\}$
$=1+\sum_{N\leq Q}\#\{(A,C)\in\mathbb{Z}^{2}:A^{2}+DB^{2}=C^{2},\gcd(A,C)=1,\ \mathrm{and}\ |C|\leq Q/N\}$
(9) $=1+\sum_{N\leq Q}\left(\frac{2^{J}}{\pi\sqrt{D}}\frac{Q}{N}+O\left(\sqrt{\frac{Q}{N}}\log\frac{Q}{N}\right)\right)$
(10) $=1+\frac{2^{J}}{\pi\sqrt{D}}Q\sum_{N\leq Q}\frac{1}{N}+\sum_{N\leq Q}O\left(\sqrt{\frac{Q}{N}}\log\frac{Q}{N}\right),$
where (9) follows from applying Theorem 3.1, and $J$ denotes the numerator of the constant appearing in Theorem 3.1 (note that the value of $J$ changes depending on whether $D$ is even or odd).
First we handle the main term in (10). By partial summation (see, for example, [HW68, Theorem 422]), we have $\sum_{N\leq Q}\frac{1}{N}=\log Q+\gamma+O(\frac{1}{Q})$, where $\gamma=0.57721$... is the Euler-Mascheroni constant. This can be simplified to $\sum_{N\leq Q}\frac{1}{N}=\log Q+O(1).$
Next, we handle the error term in (10). From the last line in [FRI77, Section 2], we have
$\sum_{N\leq Q}\frac{1}{\sqrt{N}}\log\frac{Q}{N}=O(\sqrt{Q}).$
Thus, we can bound the error term in (10) as follows:
$\sum_{N\leq Q}\sqrt{\frac{Q}{N}}\log\frac{Q}{N}=\sqrt{Q}\sum_{N\leq Q}\frac{1}{\sqrt{N}}\log\frac{Q}{N}=\sqrt{Q}\cdot O(\sqrt{Q})=O(Q).$
Therefore, we can simplify (10) to $\frac{2^{J}}{\pi\sqrt{D}}Q\log Q+O(Q).$ By plugging in the values of $J$ that appear in Theorem 3.1, we obtain the following count for all admissible integral solutions $(A,C)$:
(11) $\begin{array}[]{ll}\frac{2^{t+2m}}{\pi\sqrt{D}}Q\log Q+O(\sqrt{Q}\log Q)&\text{if $D$ is odd}\\ \frac{3\cdot 2^{t+2m-2}}{\pi\sqrt{D}}Q\log Q+O(\sqrt{Q}\log Q)&\text{if $D$ is even.}\end{array}$
@@
$\mathcal{F}^{red}_{m,D}(Q)\leq\sum_{i=1}^{m-1}\iota_{0}(i)\cdot\#\mathcal{F}^{\prime}_{m-i}(Q)\leq\sum_{i=1}^{m-1}\iota_{0}(i)\cdot\frac{\kappa(i,D)}{\pi\sqrt{D}}Q^{m-i-1}\log Q\leq\iota(m,D)Q^{m-2}\log Q$
@@ 5. Counting lengths of closed geodesics in commensurability classes
The rank of $q$ is $n+1$, which is also the degree of $p_{T}\in\mathbb{Z}[x]$. The matrix powers $T,T^{2},\dots,T^{n}$ all have entries in $\mathbb{Q}$. Let $\ell$ denote the least common multiple of the denominators appearing in the entries of these $n$ rational matrices. By the Cayley-Hamilton Theorem, $T$ satisfies its characteristic polynomial, so for any $k\geq 0$, we have
$T^{k+n+1}=-\left(T^{k}+c_{1}T^{k+1}+\cdots+c_{r-1}T^{k+n}\right).$
The subgroup of $\mathbb{Q}^{n+1}$ generated by the $T^{k}\mathbb{Z}^{n+1}$ for all $k\geq 0$ lies between the subgroups $\mathbb{Z}^{n+1}$ and $(\frac{1}{\ell}\mathbb{Z})^{n+1}$ and is thus itself a free abelian group of rank $n+1$. We can therefore choose a change of basis matrix $g\in\mathrm{GL}(n+1,\mathbb{Q})$ such that $g^{-1}Tg$ has entries in $\mathbb{Z}$.
@@
The Salem numbers which are realizable in the commensurability class of $\mathrm{SO}(q,\mathbb{Z})$ have even degree $2m\leq n+1$. Let $\mathcal{R}_{q}(Q)$ denote the set of realizable Salem numbers with $\lambda<Q$. By [MON13], the reduced determinant of a rational quadratic form of even rank is a commensurability invariant. By [BAY15, Corollary 9.3], a necessary condition for the existence of an isometry in $\mathrm{O}(q,\mathbb{Q})$ with minimal polynomial $f$ is that $\det(q)\equiv f(1)f(-1)$ in $\mathbb{Q}^{\times}/\mathbb{Q}^{\times 2}$. Since $f(0)=1$, such an isometry will be in $\mathrm{SO}(q,\mathbb{Q})$, and hence by Lemma 5.1, it will be in the commensurability class of $\mathrm{SO}(q,\mathbb{Z})$. Therefore we have
@@
We can now deduce C and D.
Proof of C and D.
Suppose first that $\Gamma$ is classical. Recall that the length of a hyperbolic element $\gamma\in\Gamma$ satisfies $\ell(\gamma)=\log\lambda$ where $\lambda$ is the unique real eigenvalue of $\gamma$ such that $\lambda>1$. By Theorem 1.1, $\lambda$ is a Salem number. Then $\ell(\gamma)\leq L$ if and only $\lambda\leq e^{L}$. C follows from B by replacing $Q$ with $e^{L}$.
Now suppose $\Gamma$ is not classical. Given a hyperbolic element $\gamma\in\Gamma$, its square $\gamma^{2}$ is in the classical finite-index subgroup $\Gamma^{(2)}\subset\Gamma$ generated by squares of elements of $\Gamma$. Then the unique real eigenvalue $\lambda>1$ of $\gamma^{2}$ is a Salem number. Here $2\ell(\gamma)=\ell(\gamma^{2})=\log\lambda$. Then $\ell(\gamma)\leq L$ if and only if $\lambda\leq e^{2L}$. D similarly follows from B by replacing $Q$ with $e^{2L}$. ∎
@@ 6. Bounds on multiplicity of the length spectrum
For arithmetic hyperbolic orbifolds, a lower bound for the average of the multiplicities can be obtained by counting Salem numbers. This study was initiated in [BLM+22] for even $n$, and in [GRE24] for odd $n$. Both of these papers used results about the so-called square-rootable Salem numbers. Let